# NUDT4MSTAR: A Large Dataset and Benchmark Towards Remote Sensing Object Recognition in the Wild

### v2

@@ 3.5 Dataset value
Robustness recognition. Our data has diverse acquisition conditions, and each image is labeled with specific information. Therefore, we can discuss in detail the robustness of the SAR recognition algorithms when the acquisition conditions (e.g., scene and sensor sensors) vary between the training and test sets. If we consider data from different acquisition conditions as domains, the training set can include multiple domains to learn domain invariant features, i.e., the domain generalization [109].
Few-shot learning. In addition to acquisition condition variations, the problem of small samples due to high costs is also a concern issue. The few-shot setting can be combined with robust recognition, and our dataset’s large number of target types can enrich the task construction of meta-learning [31, 95, 94] settings in existing few-shot learning.
@@
Incremental learning. The diversity of our dataset can support the study of domain incremental learning and class incremental learning [84]. The domain incremental learning can improve the algorithm’s robustness with a dynamic process, and class incremental learning can progressively increase the ability to recognize or reject new classes. This incremental capability is required for SAR ATR systems facing open environments.
Physical deep learning. SAR images have unique properties, such as complex phase and polarization. Our dataset provides a variety of data formats that can support recognition studies based on these unique characteristics rather than relying only on quantified SAR magnitude images. In addition, detailed metadata can also provide more gains for SAR ATR.
@@ 4 Benchmarks
We consider 7 experimental settings with 2 data formats as classification and detection benchmarks for the NUDT4MSTAR dataset in Table IV. These experimental settings include 2 SOC settings sampled from similar distributions and 5 EOC settings with obvious distribution shifts. The data formats are magnitude images in the ground range coordinate system and complex images in the slant range coordinate system.
In conclusion, we evaluate our benchmarks under 7 experimental settings and 15 recognition methods with classification and detection tasks based on magnitude images and classification tasks based on complex images. These experiments provide valuable insights and huge potential into recognizing SAR target characteristics under different imaging conditions and fine-grained classes. Our dataset is sufficient to give a new challenging benchmark to advance SAR target recognition technology. We also welcome researchers to contact us to construct richer experimental benchmarks from original data and publish them with your clear insights.
@@ 4.1 Experimental settings
TABLE IV: SOC and EOC settings of NUDT4MSTAR. The imaging conditions in SOC are similar, while EOC considers variations in a single imaging condition. Simple scenes are sandstone and bare soil, and complex scenes are urban, factory, and woodland. The separate labeling of the ground distance and slant distance images results in their numbers of annotations not being strictly corresponding. # Types.: Number of object types. Dep.: Depression angle. Azi.: Target azimuth angle. Pol.: Polarization. # Img. (Ground): Number of images in ground range. # Img. (Slant): Number of images in slant range.
@@
[…]
@@ 4.2 Results and analyses
TABLE V: Classification results. We use accuracy as a metric, i.e., the number of correctly classified samples in proportion to the total number of samples. Gray color methods use complex images, and magnitude image methods are sorted by network architecture, such as CNN and Transformer. Bolded text indicates the best result, while underlined text is the next best result.
@@
SOC-40 88.8 89.1 90.6 91.7 96.0 76.4 86.8 96.4 80.9 88.2
SOC-50 72.9 63.7 71.2 72.9 81.6 59.2 68.0 85.2 55.2 65.7
EOC-Scene 21.6 33.6 16.1 18.0 16.5 12.9 15.8 19.5 26.1 29.7
– (city) 22.7 33.7 18.1 18.9 17.9 11.6 13.9 20.4 26.1 32.7
– (factory) 20.5 34.3 13.7 16.8 14.4 15.3 17.8 18.4 25.6 30.1
– (woodland) 6.8 27.6 4.2 5.4 4.2 1.97 3.8 2.7 20.2 26.3
EOC-Depression 33.2 32.9 33.9 37.2 43.1 30.4 31.4 39.9 21.8 28.3
– (30) 52.4 54.0 52.2 55.9 63.2 43.7 47.3 58.1 39.2 46.5
– (45) 33.5 32.7 34.2 38.3 45.1 31.0 32.0 41.2 19.2 25.9
– (60) 14.0 12.3 15.4 17.4 21.3 16.7 15.0 20.5 8.1 12.7
EOC-Azimuth 15.7 16.4 14.9 16.5 21.1 29.0 22.8 26.4 19.0 22.5
– (60) 20.7 19.4 18.0 22.1 26.2 34.4 27.2 28.9 26.8 31.0
@@
Figure 14: Transfer learning results. We report the 5-shot fine-tuning performance of transfer learning on three SAR target classification datasets after NUDT4MSTAR training. The results show that our dataset has a good effect on ground target datasets such as vehicles and aircraft, but the effect is insignificant for sea targets.
@@ 4.3 Discussion
Researchers have pursued realizing a stable and efficient SAR ATR recognition system, but previous vehicle benchmark datasets have not provided difficult challenges. Our results show that current state-of-the-art methods still face many robustness and small sample difficulties. Robust feature extraction under different imaging conditions and scenes is still a worthy research problem. It is also important to efficiently use a small number of labeled samples and mine complex and polarization information. Moreover, there is no single method that can handle different perturbation conditions, demonstrating the complexity of stable SAR ATR. Of course, building a large-scale dataset and advancing SAR ATR technology is a long-term process. In a data-driven approach dominant in the 2020s, a large-scale dataset will help to extract stabilizing features and be effective in many other datasets and tasks. Based on the results, we discuss several potential directions for the future development of ATR systems.
Robustness under different conditions. As Table V and  VI show, the variations in imaging conditions present a significant challenge for recognition. This variability in the domain distribution requires the extraction of stable invariant target features under different conditions.
Efficient utilization of samples. Due to the difficulty of acquiring and labeling SAR images, it is unlikely that the sample size will reach the same size as the visible light dataset. Therefore, it is important to improve learning efficiency by utilizing a small number of samples in Fig 13.
Incorporation with multi-dimensional information. Using magnitude images alone loses many SAR image attributes, such as complex information that can effectively distinguish between targets and clutter in Table V. Therefore, auxiliary information such as complex, polarization, and angle can be introduced in SAR ATR, but increasing the dimensionality of the information may increase the instability and complexity of the features.
Fast incremental learning capability. Because of the open and dynamic world, real ATR systems face constantly changing imaging conditions and target types. They need to effectively recognize anomalies and update their capabilities in response to variations. This data provides a rich set of target types and imaging conditions that can be used to explore this possibility.
@@ 5 Conclusion
A large-scale fine-grained SAR vehicle dataset and benchmark, named NUDT4MSTAR, has been established in this paper. It contains various target types, scene variations, and imaging conditions and is 10 times the size of previous datasets of the same type. Our benchmark of 7 experiment settings and 15 methods result show that stable and efficient SAR ATR recognition method are still challengeable. It will advance diverse issue explorations and SAR ATR techniques with large-scale data. In the future, we will continue to contribute to the SAR target datasets, and we welcome researchers to contact us to jointly promote the progress of the SAR ATR field.

### v3

@@ 3.5 Dataset value
Robustness recognition. Our dataset encompasses diverse acquisition scenarios, with each image accompanied by exhaustive acquisition metadata. Therefore, we can discuss in detail the robustness of the SAR recognition algorithms when the acquisition conditions (e.g., , scene and sensor sensors) vary between the training and test sets. If we consider data from different acquisition conditions as domains, the training set can include multiple domains to learn domain invariant features, i.e., the domain generalization [100].
Few-shot learning. Besides acquisition condition variation, the inherent challenge of limited training samples due to high SAR collection costs also presents critical research opportunities. The few-shot setting can be combined with robust recognition, and our dataset’s large number of target types can enrich the task construction of meta-learning [101, 102, 103, 104] settings in existing few-shot learning.
@@
Incremental learning. The diversity of our dataset supports investigations into domain incremental learning and class incremental learning [107]. The domain incremental learning can improve the algorithm’s robustness with a dynamic process, and class incremental learning can progressively increase the ability to recognize or reject new classes. These incremental capacities are essential for SAR ATR in open environments.
Physical deep learning. SAR images have unique properties, such as complex phase and polarization. Our dataset provides multi-format data to facilitate recognition studies leveraging these intrinsic attributes instead of only relying on quantized SAR magnitude images. In addition, detailed metadata can also provide more gains for SAR ATR.
Generative Models. Beyond recognition tasks, our dataset enables controllable generation of target samples under diverse imaging conditions [2], as well as estimation of target parameters across varying acquisition scenarios.
@@ 4 Benchmarks
We consider 7 experimental settings with 2 data formats as classification and detection benchmarks from this dataset in Table IV, named ATRBench. These experimental settings include 2 SOC settings sampled from similar distributions and 5 EOC settings with obvious distribution shifts. The data formats consist of magnitude images in the ground range coordinate system and complex images in the slant range coordinate system.
In conclusion, we evaluate our benchmarks under 7 experimental settings and 15 recognition methods with classification and detection tasks based on magnitude images and classification tasks based on complex images. These experiments provide valuable insights and substantial potential into recognizing SAR target characteristics under different imaging conditions and fine-grained classes. Our dataset is sufficient to provide a new, challenging benchmark for advancing SAR target recognition technology. We also welcome researchers to contact us to construct richer experimental benchmarks from original data.
@@ 4.1 Experimental settings
TABLE IV: SOC and EOC settings of ATRNet-STAR. The imaging conditions in SOC are similar, while EOC considers variations in a single imaging condition. Simple scenes are sandstone and bare soil, and complex scenes are urban, factory, and woodland. The separate labeling of the ground distance and slant distance images results in their numbers of annotations not being strictly corresponding. # Types.: Number of object types. Dep.: Depression angle. Azi.: Target azimuth angle. Pol.: Polarization. # Img. (Ground): Number of images in ground range. # Img. (Slant): Number of images in slant range.
@@
SOC-40 train 40 all 15, 30, 45, 60 0$\sim$360 X & Ku quad 68,091 67,780
Test 40 all 15, 30, 45, 60 0$\sim$360 X & Ku quad 29,284 29,169
SOC-50 train 50 all 15, 17, 30, 45, 60 0$\sim$360 X & Ku quad 18,071 18,071
test 50 all 15, 30, 45, 60 0$\sim$360 X & Ku quad 17,603 17,613
EOC-Scene train 40 simple 15, 30, 45, 60 0$\sim$360 X & Ku quad 19,584 19,584
test 40 difficult 15, 30, 45, 60 0$\sim$360 X & Ku quad 77,791 77,365
EOC-Depression train 40 all 15 0$\sim$360 X & Ku quad 24,361 22,206
test 40 all 30, 45, 60 0$\sim$360 X & Ku quad 73,014 74,743
EOC-Azimuth train 40 all 15, 30, 45, 60 0$\sim$60 X & Ku quad 18,636 18,592
test 40 all 15, 30, 45, 60 60$\sim$360 X & Ku quad 78,739 78,357
EOC-Band train 40 except city 15, 30, 45, 60 0$\sim$360 X quad 27,711 27,653
test 40 except city 15, 30, 45, 60 0$\sim$360 Ku quad 27,763 27,732
EOC-Polarization train 40 all 15, 30, 45, 60 0$\sim$360 X & Ku HH 24,361 24,246
test 40 all 15, 30, 45, 60 0$\sim$360 X & Ku other 73,014 72,703
SOC and EOC settings. SOC settings are those where the training and test sets have similar imaging conditions, and we have SOC-40 and SOC-50 settings. SOC-40 is created from a random data sample with a 7:3 training-to-test ratio. SOC-50 randomly selects data with a similar amount of MSTAR ten classes, and we combine our dataset with MSTAR. EOC settings are those where there is a significant domain shift between the training set and the test set, such as variation in imaging conditions and target state. For EOC-Scene, we use simple scenes (sandstone and bare soil) as the training set and complex scenes (urban, factory, and woodland) with occlusion as the test set. EOC-Depression is the training set with a 15 depression angle and the test set with 30, 45 and 60. EOC-Azimuth means that training has limited target azimuth angles within 0$\sim$60, and the test has other angles. Imaging angle and geometry variation can change target signatures and background clutter. EOC-band and EOC-polarization are designed to test the effect of different bands and change from HH-polarization to other polarizations on recognition, respectively. We do not consider some experimental settings that added simulation perturbations, such as random/Gaussian noise and simulation occlusions.
@@ 4.2 Results and analyses
[…]
@@
SOC-40 $88.8$ $89.1$ $90.6$ $91.7$ $\underline{96.0}$ $76.4$ $86.8$ $\mathbf{96.4}$ $80.9$ $88.2$
SOC-50 $72.9$ $63.7$ $71.2$ $72.9$ $\underline{81.6}$ $59.2$ $68.0$ $\mathbf{85.2}$ $55.2$ $65.7$
EOC-Scene $21.6$ $\mathbf{33.6}$ $16.1$ $18.0$ $16.5$ $12.9$ $15.8$ $19.5$ $26.1$ $\underline{29.7}$
– (city) $22.7$ $\mathbf{33.7}$ $18.1$ $18.9$ $17.9$ $11.6$ $13.9$ $20.4$ $26.1$ $\underline{32.7}$
– (factory) $20.5$ $\mathbf{34.3}$ $13.7$ $16.8$ $14.4$ $15.3$ $17.8$ $18.4$ $25.6$ $\underline{30.1}$
– (woodland) $6.8$ $\mathbf{27.6}$ $4.2$ $5.4$ $4.2$ $1.97$ $3.8$ $2.7$ $20.2$ $\underline{26.3}$
EOC-Depression $33.2$ $32.9$ $33.9$ $37.2$ $\mathbf{43.1}$ $30.4$ $31.4$ $\underline{39.9}$ $21.8$ $28.3$
– (30) $52.4$ $54.0$ $52.2$ $55.9$ $\mathbf{63.2}$ $43.7$ $47.3$ $\underline{58.1}$ $39.2$ $46.5$
– (45) $33.5$ $32.7$ $34.2$ $38.3$ $\mathbf{45.1}$ $31.0$ $32.0$ $\underline{41.2}$ $19.2$ $25.9$
– (60) $14.0$ $12.3$ $15.4$ $17.4$ $\mathbf{21.3}$ $16.7$ $15.0$ $\underline{20.5}$ $8.1$ $12.7$
EOC-Azimuth 15.7 16.4 14.9 16.5 21.1 $\mathbf{29.0}$ 22.8 26.4 19.0 22.5
– (60) 20.7 19.4 18.0 22.1 26.2 $\mathbf{34.4}$ 27.2 28.9 26.8 31.0
@@
Figure 14: Transfer learning results. We report the 5-shot fine-tuning performance of transfer learning on three SAR target classification datasets after ATRNet-STAR training. The results show that our dataset has a good effect on ground target datasets such as vehicles and aircraft, but the effect is insignificant for sea targets.
@@ 4.3 Discussion
Researchers have endeavored to realize a stable and efficient SAR ATR recognition system, but previous vehicle benchmark datasets have not provided sufficiently challenging conditions. Our results show that representative methods encounter significant challenges, such as imaging angle and scene variation, as well as few-shot settings. Robust feature extraction under different varying imaging conditions remains a critical research problem. It is also crucial to efficiently utilize a small number of labeled samples and mine complex and polarization information. Although the pre-trained foundation model SARATR-X is effective in most cases compared to other methods, no single method can effectively handle diverse perturbation conditions, highlighting the complexity of SAR ATR. In a data-driven approach dominant in the 2020s, our transfer learning results show that a large-scale dataset will help to extract stabilizing features and enhance performance across multiple datasets and tasks. Based on the results, we discuss several potential directions for the future development of ATR systems.
Robustness under different conditions. As shown in Table V and  VI, the variations in imaging conditions present a significant challenge for recognition. This variability in the domain distribution requires the extraction of stable invariant target features under different conditions.
Efficient utilization of samples. Given the challenges associated with acquiring and labeling SAR images, it is unlikely that the sample size will reach the same size as the visible light dataset. Therefore, it is important to improve learning efficiency by utilizing a small number of samples in Fig 13.
Incorporation with multi-dimensional information. Table V show that complex information can help distinguish between targets and clutter under scene variations. Auxiliary information such as complex and angle can be introduced in SAR ATR, but increasing the dimensionality of the information may increase the instability and complexity.
Fast incremental learning capability. Because of the open real-world ATR systems face constantly changing imaging conditions and target types. They need to effectively recognize anomalies and update their capabilities in response to variations. This dataset provides a rich set of target types and imaging conditions that can be used to explore this possibility.
@@ 5 Conclusion
A large-scale fine-grained SAR vehicle dataset and benchmark, named ATRNet-STAR, has been established in this paper. It contains various target types, scene variations, and imaging conditions and is 10 times the size of previous datasets of the same type. We provide a comprehensive benchmark of 7 experiment settings and 15 methods. It will facilitate diverse explorations of issues and promote SAR ATR techniques through large-scale data. In the future, we will continue to contribute to the SAR target datasets, and we welcome researchers to contact us to jointly promote the progress of the SAR ATR field.